# AI-Informed Model Analogs for Subseasonal-to-Seasonal Prediction

### v2

@@ Abstract
Subseasonal-to-seasonal (S2S) forecasting is crucial for public health, disaster preparedness, and agriculture, yet it remains particularly challenging to forecast and to understand the modes of predictability on this timescale. We adapt an interpretable AI-informed analog forecasting approach, previously used for longer timescales, to improve S2S analog prediction and understanding of its climate drivers. Using an artificial neural network, we learn a mask of weights to optimize analog selection and showcase its versatility across two prediction tasks: 1) regional continuous prediction of Month 1 midwestern U.S. summer temperatures and 2) classification of Month 1-2 North Atlantic wintertime upper atmospheric winds. The AI-informed analogs outperform traditional analog forecasting approaches, as well as climatology and persistence baselines, for deterministic and probabilistic skill metrics on both climate model and reanalysis data; moreover, this skill gap grows for extreme predictions. Moreover, by using an interpretable-AI framework, we analyze the learned masks of weights to understand the underlying physical processes and find skin temperature and the Northern Hemisphere to be important predictors of North Atlantic wintertime upper atmospheric winds.
: Machine Learning: Earth
@@ 1 Introduction
Forecasting on S2S timescales, typically defined as 2 weeks to $\sim$2 months, is vital for public health, disaster preparedness, agriculture, and energy/water management [White:2017aa]. Despite the clear benefits of skillful predictions on these timescales, S2S forecasting remains especially difficult. Often referred to as a ‘predictability desert’ [Robertson:2018ab, Chen:2024aa], S2S forecasts cannot solely rely on the initial atmospheric conditions, as is often done in short-term numerical weather prediction, or on the slow-varying boundary conditions that underpin climate outlooks [Robertson:2018ab, Vitart:2018aa]. Instead, forecasters must integrate information from initial conditions, boundary conditions, and S2S modes of variability, like the Madden Julian Oscillation (MJO) [Zhang:2013aa], to produce skillful predictions [Vitart:2018aa]. Still, on S2S timescales, the strength of these sources of predictability and their teleconnections remain unclear [Merryfield:2020aa, Vitart:2018aa] and skill, e.g. accuracy of summertime surface temperature prediction in North America, remains relatively low [Breeden:2022aa, Pegion:2019aa].
A variety of tools have been used to approach the S2S forecasting challenge. Dynamical models have slowly but steadily improved S2S forecast skill [Peng:2023aa] and data-driven approaches, like fully-AI models, can now forecast phenomena such as the North Atlantic Oscillation (NAO) and MJO at S2S lead times [ling2024fengwuw2sdeeplearningmodel, Chen:2024aa] with similar skill to dynamical models. To further improve forecasts, there has recently been a renewed focus on pinpointing climate states that represent times of enhanced predictability [[, e.g.,]]Mariotti:2020aa,Mayer:2021aa, Albers:2019aa. Identifying these ‘windows of opportunity’ is a potential approach to improve skill on S2S timescales by allowing forecasters to know when forecast uncertainty is high or when they can leverage these times of enhanced predictability for more accurate forecasts [Mariotti:2020aa].
Here, we tackle S2S prediction by combining a variety of these methodologies and employing an AI-informed model analog forecasting approach. Analog forecasting rests on the premise that climate states with similar initial conditions tend to evolve in a consistent manner [[, e.g.,]]Lorenz:1969aa,Zhao:2016aa. By identifying past states resembling current conditions, their subsequent evolution can offer plausible trajectories for future conditions. For a variety of forecasts, from the tropics to the northern high latitudes, analog forecasting has been shown to rival the skill of global climate models [Lou:2023aa, Ding:2019aa, Walsh:2021aa] all while offering several key advantages. Unlike fully-AI models, analog forecasting is intuitive, interpretable, and can uphold physical laws [Rader:2023aa, Ding:2018aa]; moreover, compared to global dynamic climate models, analog forecasting is highly computationally efficient [Ding:2019aa].
@@
To address this need for a large analog library, we turn to climate models, which have orders of magnitude more climate realizations than we have observational data [Ding:2018aa, McDermott:2016aa]. Yet, even with climate models, finding perfect analogs is impractical—estimates suggest over $10^{30}$ years of data would be needed to match two atmospheric flow stream patterns in just the Northern Hemisphere within observational error [VAN-DEN-DOOL:1994aa]. Hence, determining the conditions that make a climate state an adequately close analog, rather than a perfect one, is crucial. For example, [Ding:2018aa] use regional matching to identify close analogs for seasonal tropical Indo-Pacific Ocean prediction; [esd-13-1437-2022] use global matching for multi-decadal global predictions; and [Wu:2023aa] use area-specific matching for annual-to-multi-year Pacific Decadal Oscillation prediction. These methods for selecting analogs have been shown to work for certain problems, although they demand either a huge library of analogs (as in global matching) or depend on prior knowledge of physical drivers and teleconnections (as in regional or area-specific matching).
Here, we explore an alternative, AI-based spatial weighting approach originally introduced by [Rader:2023aa]. We train a neural network to output a mask of weights that highlights where it is most important for initial conditions to match, such that two states will evolve similarly. Using a learned set of weights to find optimal analogs reduces reliance on prior knowledge and enables investigation of which regions and variables are most essential for two climate states to follow similar future trajectories. This method of optimized analog forecasting was first successfully applied to annual-to-decadal sea surface temperature prediction [Rader:2023aa] and has since been extended to multi-year-to-decadal 2-meter temperatures [fernandez2025multiyeartodecadaltemperaturepredictionusing] and seasonal-annual El Niño-Southern Oscillation (ENSO) predictions [toride2024usingdeeplearningidentify].
Here, we show that this AI-based analog forecasting approach can achieve skill beyond traditional analog methods on S2S timescales while maintaining interpretability and computational efficiency. We highlight the benefits of using AI-based analogs across two prediction tasks: 1) regional continuous prediction of Month 1 midwestern U.S. summer temperatures and 2) classification of Month 1-2 North Atlantic wintertime upper atmospheric winds. Through these predictions tasks we show the AI-based analog approach outperforms traditional analog forecasting approaches, climatology, and persistence on reanalysis data on S2S timescales, exhibiting especially strong performance for extreme temperature prediction. Moreover, via this interpretable AI-forecasting framework we analyze the learned masks of weights to better understand the S2S sources of predictability that underpin this improvement.
@@ 2.1 Prediction Tasks
We demonstrate the skill of our analog forecasting approach for the prediction tasks described in Table 1, opting for a varied pair of examples to test the generalizability of the method across different S2S prediction problems. We apply the AI-informed analog approach to both classification and continuous prediction tasks, to different regions, seasons, variables, and lead times. These prediction tasks each have a unique learned mask of weights to optimize the choice of analogs. While we focus on monthly prediction, we also find skill over traditional analog methods for week 3-4 prediction in Task #3, included in the Supplemental Section S1.
@@ 2.3 Data
We predict both U250 and skin temperature, relying on output from the Community Earth System Model 2–Large Ensemble (CESM2-LE) [Danabasoglu:2020aa] in order to have a sufficiently large analog library and to learn the weighted mask. As we will show, the AI-informed analog approach produces skillful predictions when evaluated on both CESM2-LE data, in a perfect-model framework, and on ECMWF Reanalysis v5 (ERA5) data [Hersbach:2020aa]. We make use of monthly-mean fields that is resolved at .25° x .24° (natively for CESM2-LE data, and bilinearly interpolated for ERA5 data). For all data sources, we convert the data to anomalies about the climatological seasonal cycle and then to standard deviations across the subsetted season at each grid point. However, between data sources, we handle the anthropogenic effects of climate change slightly differently, as will be discussed next.
@@ 2.3.1 CESM2-LE Data
We use monthly CESM2-LE data from 1850-2100 that employs CMIP6 historical and SSP3-7.0 future radiative forcing scenarios [Simpson:2023aa]. We take all 100 members to calculate the ensemble mean, which we subtract from each individual member to both remove the effects of anthropogenic climate change and to convert the data to anomalies from the seasonal cycle. To increase speed and reduce memory load, we then use only a third of the members for training and the analog library. These members are divided between the analog library and SOIs, with fields from 19 members composing the library and fields from 14 members serving as the SOIs (see Table S2 for member details). We partition the SOIs with a 10/2/2 member split for training, validation, and testing respectively.
@@ 2.3.2 ERA5 Data
We use ERA5 data from January 1940 to July 2024. We fit and subtract a third-order polynomial at each grid point and each calendar month to define detrended anomalies from the seasonal cycle. The ERA5 data acts as a second test set to evaluate skill on observations.
@@ 2.4 Artificial Neural Network to Learn Mask of Weights
The goal of the artificial neural network is to optimize the premise of analog forecasting; namely, that smaller differences in initial conditions lead to smaller differences in future conditions. To optimize this relationship between initial and future conditions, we task the network with predicting how similarly two states will evolve in the target region given their initial conditions. We employ a network that is similar to that of [Rader:2023aa], with minor modifications to its final layers. During each forward pass, the network, depicted in Figure 2, takes two maps as input and uses these initial conditions to predict the similarity of their future states. The SOI map is from the training set and the analog map is randomly selected from the analog library. These maps are both multiplied by a grid of learnable weights (i.e., the mask) of the same size as the inputs, resulting in two weighted maps. The mask is restricted to have a mean of 1 across all weights, such that during training weight is moved between different areas of the globe, but conserved. The MSE between these two weighted maps is calculated, representing the effective similarity of the two initial states for this prediction task, and is passed through a single linear scaling layer. The output of this layer represents the network’s prediction of the MSE between the two maps in the target region after they have evolved (i.e., after the desired lead time). Loss is computed as the MSE between the predicted difference of the targets and the true difference of the targets. Hence, through this weighting process, the network learns a mask that aligns the MSE between two states’ inputs to their MSE after evolution in the target region. This process is repeated for each SOI in the training set. Details of the network setup and hyperparameters can be found in Table S1.
Our network deviates from that of [Rader:2023aa], in that we use a single linear layer instead of multiple dense layers at the end of the network. We restrict the linear layer’s weight to be $\geq 0$ to ensure a monotonically increasing relationship between the MSE of the maps’ weighted inputs and the predicted MSE of the maps’ target regions. This better matches our process for selecting analogs as described in Section 2.2, as we expect that two maps with a smaller weighted MSE will also evolve to have a smaller MSE between their targets. This switch to a single linear layer resulted in a negligible change in skill, but increased network parsimony and training speed.
@@ 2.5.1 Baselines
We include a regional and a global analog baseline in addition to persistence, climatological, and random baselines to evaluate the relative skill of the learned mask approach. To create a global baseline, we select analogs by matching conditions over the entire globe (equivalent to a weighted mask of 1s everywhere). We create a regional baseline by selecting analogs via matching conditions only in the target region (equivalent to a weighted mask of 1s in the target region and 0s everywhere else). The 90th percentile random baseline is formed by repeating the prediction for all SOIs using random analogs 100 times and selecting the 90th percentile of best predictions.
@@ 3 Results
[…]
@@
The learned mask modestly outperforms all baselines (Figure 4), with MAE skill increases of 17% and 51% and CRPS skill increases of 5% and 48% tested on CESM2-LE and ERA5, respectively. All skill scores peak at 50 analogs for both CESM2-LE and ERA5 data, except for CESM2-LE CRPS, which peaks at 100 analogs. This overall improvement in temperature forecasting on S2S timescales is important, however, better prediction of extreme temperatures in particular has an outsized impact on enhancing agricultural production, public health, and energy management [Domeisen:2022aa]. Thus, we focus on assessing the AI-based analog’s ability to predict extreme temperatures. Here, we utilize a discard plot, in which we progressively discard samples with lower extremity to visualize how MAE skill changes for more extreme samples. We denote extremity simply as the absolute value of the prediction, i.e. a measure of how far from climatology the prediction is.
@@
In Figure 5, we show the discard plot with ERA5 data using an ensemble of 50 analogs. The AI-based analogs exhibits a marked increase in skill for samples with more extreme predictions. As our skill score is defined relative to climatology, it may be unsurprising that the AI-informed analogs would have lower relative error on more extreme events. However, this is not the case for the regional baseline, where there is only a slight increase in skill for the most extreme samples. This analysis highlights how the skill gap between AI-informed analogs and traditionally-selected analogs widens for more extreme temperature events. Moreover, as we use predicted extremity to discard samples, this information is available a priori, allowing forecasters to better understand when the analog ensemble is likely to perform best and building trust in its more extreme predictions. This behavior also holds for CESM2-LE data (Figure S6).
@@ 3.2.1 Physical Interpretation of the Learned Mask
Using this interpretable AI-informed analog approach, we can analyze the learned mask to better understand the physical drivers behind analog predictions. Here we discuss three areas of high weight in the mask, shown in Figure 6: 1) SSTs in the eastern tropical Pacific, 2) SSTs in the Phillipine Sea, and 3) U250 in the Northern Hemisphere. Area 1 resembles the canonical ENSO region. We find this reflected in the analog selection, with selected analogs generally have more similar ENSO states than a random selection (Figure S10), as measured by Niño-3.4 indices [AVerifiedEstimationoftheElNioIndexNio34since1877]. This is in line with work that finds ENSO impacts the North Atlantic region through its influence on the North Atlantic Oscillation (NAO)—the most prominent pattern of atmospheric variability in the region [ENSO_NAO, NAO]. Area 2 resembles a prominent Rossby wave source region, which is a known source of MJO teleconnections to the North Atlantic [MJO_NAO_teleconnections, TheConsistencyofMJOTeleconnectionPatternsAnExplanationUsingLinearRossbyWaveTheory]. Lastly, Area 3 highlights the Pacific jet exit region of the subtropical jet stream. The region of high weight over the Eastern Pacific and Mexico aligns with the mean position of the subtropical jet, and both its strength and the two North–South shifted branches, shown in smaller green boxes in Figure 6, have been found to vary with the NAO state [JET_NAO].
@@
We focus on a 400 analog ensemble, as this is the number of analogs for which skill peaks (Figure 9), although the general trends remain similar across ensemble size (Figure S11). We find a slight improvement in skill when we increase mask sparsity by thresholding or by introducing constrained inverse $L_{2}$ regularization. This increase in skill with a sparser map is consistent with [Rader:2023aa], who found a slight improvement in skill for multi-year predictions using a $\sim 95\%$ percentile threshold. However, when we test increasing the sparsity for shorter timescales (e.g. Task #3), we find minimal change in observation skill and a slight decrease in probabilistic model skill (Figure S9). Considering field ablation, temperature appears to be the more important of the two fields, although ablating either the temperature field or the U250 field results in a significant decrease in skill, highlighting the importance of both for identifying skillful analogs. While all ablation methods besides increasing sparsity decreases skill, ablating the Northern Hemisphere, both fields, and ocean temperatures result in the largest drop in skill.

### v3

# AI-Informed Model-Analogs for Understanding Subseasonal-to-Seasonal Jet Stream and North American Temperature Predictability

@@ Abstract
Subseasonal-to-seasonal (S2S) forecasting is crucial for public health, disaster preparedness, and agriculture, yet both forecasting and diagnosing sources of potential forecast skill on this timescale remains particularly challenging. We adapt an interpretable AI-informed analog forecasting approach, previously used for longer timescales, to improve S2S model-analog prediction and understanding of its climate drivers. Using an artificial neural network, we learn a mask of weights to optimize analog selection and showcase its versatility across two prediction tasks: 1) regional continuous prediction of Month 1 midwestern U.S. summer temperatures and 2) classification of Month 1-2 North Atlantic wintertime upper atmospheric winds. The AI-informed analogs outperform traditional model-analog forecasting approaches, as well as climatology and persistence baselines, for deterministic and probabilistic skill metrics on both climate model and reanalysis data; moreover, this skill gap grows for extreme predictions. Moreover, our interpretable-AI framework allows analysis of learned masks of weights, yielding improved understanding of the role of underlying physical processes upon predictability. We find skin temperature and the Northern Hemisphere to be more important predictors of North Atlantic wintertime upper atmospheric winds than upper atmospheric winds 1-2 months prior and the Southern Hemisphere, respectively.
@@ 1 Introduction
Forecasting on S2S timescales, typically defined as 2 weeks to $\sim$2 months, is vital for public health, disaster preparedness, agriculture, and energy/water management [34]. Despite the clear benefits of skillful predictions on these timescales, S2S forecasting remains especially difficult. Often referred to as a ‘predictability desert’ [25, 5], S2S forecasts cannot solely rely on the initial atmospheric conditions, as is often done in short-term numerical weather prediction, or on the slow-varying boundary conditions that underpin climate outlooks [25, 31]. Instead, forecasters must integrate information from initial conditions, boundary conditions, and S2S modes of variability, like the Madden Julian Oscillation (MJO) [38], to produce skillful predictions [31]. Still, on S2S timescales, the strength of these sources of predictability and their teleconnections remain unclear [21, 31] and skill, e.g. accuracy of summertime surface temperature prediction in North America, remains relatively low [3, 22].
A variety of tools have been used to approach the S2S forecasting challenge. Dynamical models have slowly but steadily improved S2S forecast skill [23] and data-driven approaches, like fully-AI models, can now forecast phenomena such as the North Atlantic Oscillation (NAO) and MJO at S2S lead times [16, 5] with similar skill to dynamical models. To further improve forecasts, there has recently been a renewed focus on pinpointing climate states that represent times of enhanced predictability [[, e.g.,]]Mariotti:2020aa,Mayer:2021aa, Albers:2019aa. Identifying these ‘windows of opportunity’ is a potential approach to improve skill on S2S timescales by allowing forecasters to know when forecast uncertainty is high or when they can leverage these times of enhanced predictability for more accurate forecasts [19].
Here, we tackle S2S prediction by combining a variety of these methodologies and employing an AI-informed model-analog forecasting approach. Analog forecasting rests on the premise that climate states with similar initial conditions tend to evolve in a consistent manner [[, e.g.,]]Lorenz:1969aa,Zhao:2016aa. By identifying past states resembling current conditions, their subsequent evolution can offer plausible trajectories for future conditions. For a variety of forecasts, from the tropics to the northern high latitudes, analog forecasting has been shown to rival the skill of global climate models [17, 8, 32] all while offering several key advantages. Unlike fully-AI models, analog forecasting is intuitive, interpretable, and can uphold physical laws [24, 7]; moreover, compared to global dynamic climate models, analog forecasting is highly computationally efficient [8].
@@
To address this need for a large analog library, we turn to climate models, which have orders of magnitude more climate realizations than we have observational data [7, 35]. Yet, even with climate models, finding perfect analogs is impractical—estimates suggest over $10^{30}$ years of data would be needed to match two atmospheric flow stream patterns in just the Northern Hemisphere within observational error [30]. Hence, determining the conditions that make a climate state an adequately close analog, rather than a perfect one, is crucial. For example, [7] use regional matching to identify close analogs for seasonal tropical Indo-Pacific Ocean prediction; [18] use global matching for multi-decadal global predictions; and [35] use area-specific matching for annual-to-multi-year Pacific Decadal Oscillation prediction. These methods for selecting analogs have been shown to work for certain problems, although they demand either a huge library of analogs (as in global matching) or depend on prior knowledge of physical drivers and teleconnections (as in regional or area-specific matching).
Here, we explore an alternative, AI-based spatial weighting approach originally introduced by [24]. We train a neural network to output a mask of weights that highlights where it is most important for initial conditions to match, such that two states will evolve similarly. Using a learned set of weights to find optimal analogs reduces reliance on prior knowledge and enables investigation of which regions and variables are most essential for two climate states to follow similar future trajectories. This method of optimized analog forecasting was first successfully applied to annual-to-decadal sea surface temperature prediction [24] and has since been extended to multi-year-to-decadal 2-meter temperatures [10] and seasonal-annual El Niño-Southern Oscillation (ENSO) predictions [28].
Here, we show that this AI-based model-analog forecasting approach can achieve skill beyond traditional model-analog methods on S2S timescales while maintaining interpretability and computational efficiency. We highlight the benefits of using AI-based analogs across two prediction tasks: 1) regional continuous prediction of Month 1 midwestern U.S. summer temperatures and 2) classification of Month 1-2 North Atlantic wintertime upper atmospheric winds. Through these predictions tasks we show the AI-based model-analog approach outperforms traditional model-analog forecasting approaches, climatology, and persistence on reanalysis data on S2S timescales, exhibiting especially strong performance for extreme temperature prediction. Moreover, via this interpretable AI-forecasting framework we analyze the learned masks of weights to better understand the S2S sources of predictability that underpin this improvement.
@@ 2.1 Prediction Tasks
We demonstrate the skill of our model-analog forecasting approach for the prediction tasks described in Table 1, opting for a varied pair of examples to test the generalizability of the method across different S2S prediction problems. We apply the AI-informed model-analog approach to both classification and continuous prediction tasks, to different regions, seasons, variables, and lead times. These prediction tasks each have a unique learned mask of weights to optimize the choice of analogs. While we focus on monthly prediction, we also find skill over traditional model-analog methods for week 3-4 prediction in Task #3, included in the Supplemental Section S1.
@@ 2.3 Data
We predict both U250 and skin temperature, relying on output from the Community Earth System Model 2–Large Ensemble (CESM2-LE) [6] in order to have a sufficiently large analog library and to learn the weighted mask. As we will show, the AI-informed model-analog approach produces skillful predictions when evaluated on both CESM2-LE data, in a perfect-model framework, and on ECMWF Reanalysis v5 (ERA5) data [11]. We make use of monthly-mean fields that are approximately resolved at 1.25° x .9° (natively for CESM2-LE data, and bilinearly interpolated for ERA5 data). For all data sources, we convert the data to anomalies about the climatological seasonal cycle and then normalize at each grid point by dividing by the standard deviation of the anomalies. However, between data sources, we handle the anthropogenic effects of climate change slightly differently, as will be discussed next.
@@ 2.3.1 CESM2-LE Data
We use monthly CESM2-LE data from 1850-2100 that employs CMIP6 historical and SSP3-7.0 future radiative forcing scenarios [27]. We take all 100 members to calculate the ensemble mean, which we subtract from each individual member to both remove the effects of anthropogenic climate change and to convert the data to anomalies from the seasonal cycle. To increase speed and reduce memory load, we then use only a third of the members for training and the analog library. These members are divided between the analog library and SOIs, with fields from 19 members composing the library and fields from 14 members serving as the SOIs (see Table S2 for member details). We partition the SOIs with a 10/2/2 member split for training, validation, and testing respectively.
@@ 2.3.2 ERA5 Data
We use ERA5 data from January 1940 to September 2024. We fit and subtract a third-order polynomial at each grid point and each calendar month to define detrended anomalies from the seasonal cycle. The ERA5 data acts as a second test set to evaluate skill on observations.
@@ 2.4 Artificial Neural Network to Learn Mask of Weights
The goal of the artificial neural network is to optimize the premise of analog forecasting; namely, that smaller differences in initial conditions lead to smaller differences in future conditions. To optimize this relationship between initial and future conditions, we task the network with predicting how similarly two states will evolve in the target region given their initial conditions. We employ a network that is similar to that of [24], with minor modifications to its final layers. During each forward pass, the network, depicted in Figure 2, takes two maps as input and uses these initial conditions to predict the similarity of their future states. The SOI map is from the training set and the analog map is randomly selected from the analog library. These maps are both multiplied by a grid of learnable weights (i.e., the mask) of the same size as the inputs, resulting in two weighted maps. The mask is restricted to have a mean of 1 across all weights, such that during training weight is moved between different areas of the globe, but conserved. The MSE between these two weighted maps is calculated, representing the effective similarity of the two initial states for this prediction task, and is passed through a single linear scaling layer. The output of this layer represents the network’s prediction of the MSE between the two maps in the target region after they have evolved (i.e., after the desired lead time). Loss is computed as the MSE between the predicted difference of the targets and the true difference of the targets. Hence, the primary objective of the network is to align through this weighting process, the network learns a mask that aligns the MSE between two states’ inputs to their MSE after evolution in the target region. This process is repeated for each SOI in the training set. Details of the network setup and hyperparameters can be found in Table S1.
Our network deviates from that of [24], in that we use a single linear layer instead of multiple dense layers at the end of the network. We restrict the linear layer’s weight to be $\geq 0$ to ensure a monotonically increasing relationship between the MSE of the maps’ weighted inputs and the predicted MSE of the maps’ target regions. This better matches our process for selecting analogs as described in Section 2.2, as we expect that two maps with a smaller weighted MSE will also evolve to have a smaller MSE between their targets. This switch to a single linear layer resulted in a negligible change in skill, but increased network parsimony and training speed.
@@ 2.5.1 Baselines
We include a regional and a global analog baseline in addition to persistence, climatological, and random baselines to evaluate the relative skill of the learned mask approach. To create a global baseline, we select analogs by matching conditions over the entire globe (equivalent to a weighted mask of 1s everywhere). We create a regional baseline by selecting analogs via matching conditions only in the target region (equivalent to a weighted mask of 1s in the target region and 0s everywhere else). The 90th percentile random baseline is formed by repeating the prediction for all SOIs using random analogs 100 times and selecting the 90th percentile of best predictions. Furthermore, we include a limited comparison of analogs in Task #1 to operational S2S models in Supplemental Section S4 as a reference point.
@@ 3 Results
We first explore how well the AI-informed model-analog forecasting approach can perform continuous prediction, by assessing monthly summer midwestern U.S. ($36^{\circ}-49^{\circ}\mathrm{N},\ 90^{\circ}-106^{\circ}\mathrm{W}$) temperatures (Task #1) with a focus on extremes. We include this focus on extreme heat prediction, as the midwestern U.S. experiences some of the highest heat index events in the country [[, e.g.,]]Romps_2022. We predict the normalized temperature anomalies each month from July through September, using the learned mask in Figure 3. The mask displays a strong emphasis on the target region, highlighting the regional importance of the central U.S. for predicting midwestern summer temperatures, and preferential weighting in the mid-latitudes of the Northern Hemisphere as well as the Maritime Continent. The Maritime Continent signal is reminiscent of the MJO, which is influential in Midwest summer climate [33], while the mid-latitude pattern resembles wave trains correlated with summertime North American heatwaves [37]. The local signal in the central U.S. itself may be a result of the strong summertime land-atmosphere and soil moisture-temperature coupling in this region [20].
@@
The learned mask modestly outperforms all baselines (Figure 4), with MAE skill increases of 17% and 51% and CRPS skill increases of 5% and 48% tested on CESM2-LE and ERA5, respectively. All skill scores peak at 50 analogs for both CESM2-LE and ERA5 data, except for CESM2-LE CRPS, which peaks at 100 analogs. This overall improvement in temperature forecasting on S2S timescales is important, however, better prediction of extreme temperatures in particular has an outsized impact on enhancing agricultural production, public health, and energy management [9]. Thus, we focus on assessing the AI-based analog’s ability to predict extreme temperatures. Here, we utilize a discard plot, in which we progressively discard samples with lower extremity to visualize how MAE skill changes for more extreme samples. We denote extremity simply as the absolute value of the prediction, i.e. a measure of how far from climatology the prediction is.
@@
In Figure 5, we show the discard plot with ERA5 data using an ensemble of 50 analogs. The AI-based model-analogs exhibits a marked increase in skill for samples with more extreme predictions. As our skill score is defined relative to climatology, it may be unsurprising that the AI-informed model-analogs would have lower relative error on more extreme events. However, this is not the case for the regional baseline, where there is only a slight increase in skill for the most extreme samples. This analysis highlights how the skill gap between AI-informed model-analogs and traditionally-selected model-analogs widens for more extreme temperature events. Moreover, as we use predicted extremity to discard samples, this information is available a priori, allowing forecasters to better understand when the analog ensemble is likely to perform best and building trust in its more extreme predictions. This behavior also holds for CESM2-LE data (Figure S6).
@@ 3.2.1 Physical Interpretation of the Learned Mask
Using this interpretable AI-informed model-analog approach, we can analyze the learned mask to better understand the physical drivers behind analog predictions. Here we discuss three areas of high weight in the mask, shown in Figure 6: 1) SSTs in the eastern tropical Pacific, 2) SSTs in the Phillipine Sea, and 3) U250 in the Northern Hemisphere. Area 1 resembles the canonical ENSO region. We find this reflected in the analog selection, with selected analogs generally have more similar ENSO states than a random selection (Figure S10), as measured by Niño-3.4 indices [4]. This is in line with work that finds ENSO impacts the North Atlantic region through its influence on the North Atlantic Oscillation (NAO)—the most prominent pattern of atmospheric variability in the region [26, 13]. Area 2 resembles a prominent Rossby wave source region, which is a known source of MJO teleconnections to the North Atlantic [15, 29]. Lastly, Area 3 highlights the Pacific jet exit region of the subtropical jet stream. The region of high weight over the Eastern Pacific and Mexico aligns with the mean position of the subtropical jet, and both its strength and the two North–South shifted branches, shown in smaller green boxes in Figure 6, have been found to vary with the NAO state [12].
@@
We focus on a 400 analog ensemble, as this is the number of analogs for which skill peaks (Figure 9), although the general trends remain similar across ensemble size (Figure S11). We find a slight improvement in skill when we increase mask sparsity by thresholding or by introducing constrained inverse $L_{2}$ regularization. This increase in skill with a sparser map is consistent with [24], who found a slight improvement in skill for multi-year predictions using a $\sim 95\%$ percentile threshold. However, when we test increasing the sparsity for shorter timescales (e.g. Task #3), we find minimal change in observation skill and a slight decrease in probabilistic model skill (Figure S9). Considering field ablation, temperature appears to be the more important of the two fields, although ablating either the temperature field or the U250 field results in a significant decrease in skill, highlighting the importance of both for identifying skillful analogs. While all ablation methods besides increasing sparsity decreases skill, ablating the Northern Hemisphere, both fields, and ocean temperatures result in the largest drop in skill.